Fermi National Laboratory
 
Operation Summary for April 13, 2012 - April 16, 2012
 

NuMI reported receiving 9.15E18 protons on target for the period from 4/9/12 to 4/16/12.

Friday April 13

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with Muon Ring personnel conducting a study, and with the NuMI, MiniBooNE, SeaQuest, and Meson FTBF T-992 experiments taking beam.

During the day shift, the NuMI beam permit tripped off at 8:44 AM for the first of two times in six hours due to a power supply’s output (E:V118) drifting low; everything reset. A Booster person began a study at 8:49 AM. The Muon Ring study ended at 8:51 PM. Operations turned off the Linac beam at 9 AM so that Linac personnel could work on a Linac RF station’s (LRF1) cavity water pump. SeaQuest personnel began an access around 9:04 AM. Linac personnel completed their work on LRF1 at 10:30 AM and Operations established reestablished beam. Muon Ring personnel began another study at 10:39 AM. A MI person began a one-hour slow extraction study at 12:41 PM. SeaQuest came out of access and resumed taking beam at 2:26 PM. The Muon Ring study ended at 2:56 PM.

During the evening shift, Operations reported at 4:33 PM that they had received the Operational Readiness Clearance (ORC) for the MTA High Pressure Cavity and began taking beam at 4:52 PM. Operations established beam to T-1005 at 7:05 PM. MTA finished taking beam at 8:13 PM. At 11:04 PM, Operations reported completing a 750 KeV area line tune. T-1005 made two accesses during this shift.

Saturday April 14

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1005.

During the midnight shift, at 12:05 AM, Operations reported completing a Booster injection tune. Operations established beam to experiment T-992 at 5:41 AM.

During the day shift, at 12:12 PM, a Booster RF station (BRF5) tripped off and would not reset; an RF person came in and had the station back online by 1:12 PM. The NuMI beam permit tripped off at 1:23 PM for the first of two times in 15 minutes due to V118; everything reset. T-992 made accesses during this shift.

During the evening shift, Operations established beam to T-1005 around 6:10 PM. At 8:30 PM, BRF11 tripped off for the first of five times in about three hours; it reset. T-1005 made an access during this shift.

Sunday April 15

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1005.

During the midnight shift, at 12:03 AM, BRF11 tripped off for the first of five times in about six hours; it reset. At 2:39 AM, Operations reported that the MI dampers and amplifiers had tripped off due to a controls problem; Operators rebooted the damper VME and turned the amplifiers back on. At 2:44 AM, a lightning strike cause two Switchyard power supplies (HP3US and HP3DS) to trip off; everything reset. At 4:54 AM, the NuMI beam permit tripped off due to V118; everything reset. Operations established beam to T-992 around 6:06 AM.

During the day shift, at 10:57 AM, BRF11 and 12 tripped off for the first of two times in about an hour; they reset.

During the evening shift, Operations established beam to T-1005 around 4:06 PM. SeaQuest made an access during this shift.

Monday April 16

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1005.

During the midnight shift, at 4:11 AM, a NuMI power supply (E:HV101) tripped off and would not reset; Operations called in Power Supply personnel. Operations halted MI beam for the HV101 repairs. Power Supply personnel reported that they had repaired two SCRs and a firing circuit; Operations reestablished MI beam at 7:40 AM. Muon Ring personnel began a study around 7:59 AM. SeaQuest accessed during this shift.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of Booster, Muon Ring, and MI personnel to conduct studies. SeaQuest continues to commission their beam line.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/16/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory